# Towards Agentic Intelligence for Materials Science

### v1

@@ 4.4 Pipeline-Centric Perspective
Figure 9: The essential role of human-in-the-loop in autonomous discovery. While depicting the cycle of agentic hypothesis generation and execution, this schematic emphasizes that human oversight remains a critical component of the pipeline. It illustrates how human experts do not just monitor for safety but actively co-improve the system, providing the high-level reasoning and validation necessary to align autonomous agents with meaningful scientific objectives.
After addressing Q3, we continue our discussion on the end-to-end pipeline-centric perspective, focusing more on the role of agentic LLMs in materials discovery. Yann LeCun has famously argued that “if intelligence is a cake, the bulk of the cake is unsupervised learning, the icing on the cake is supervised learning, and the cherry on the cake is reinforcement learning” (NeurIPS 2016), thereby downplaying the centrality of reinforcement learning relative to representation learning via large-scale pre-training. In contrast, Silver and Sutton 324 argue that intelligence should fundamentally be grounded in interaction with the environment, thereby elevating the role of reinforcement learning and other experience-driven approaches to a primary mechanism for building general intelligence. Rather than engaging directly with this broader conceptual debate, we instead focus on the concrete demands of materials science, where discovery is inherently interactive: hypotheses must be tested through simulations and (preferably) experiments, so learning through experience is less optional than it is in purely digital areas.
@@
First, a lot of current work doesn’t use backward credit assignment all the way through the pipeline. By backward credit assignment, we mean linking the results of downstream discoveries—both successes and failures—back to choices made earlier in the process, such as corpus selection, objective design, and model adaptation. This way, those parts can be changed under the same end-to-end pressure. As outlined in Sec. 2 and 3, training signals resulting from agentic successes or failures (such as unsuccessful synthesis attempts, inaccurate property predictions, and breached physical constraints) can theoretically be transmitted backward to prior stages, including pretraining data curation, domain adaptation objectives, and instruction-tuning criteria. This backward signal is necessary for the influence-driven correction discussed in Sec. 2.4, where failures in closed-loop discovery immediately tell us which parts of the pretraining corpus should be made louder, quieter, or eliminated. Without this feedback loop, agentic systems stay separate from the basic representations and assumptions built into their models. This makes it harder for them to change and get better over long periods of time.
@@ 5 Discussion & Conclusion
This survey synthesizes recent advances in AI for materials science through a pipeline‑centric lens that connects corpus curation, pretraining, domain adaptation, and instruction tuning to goal‑conditioned, agentic LLMs operating within open‑ended experimental environments. By jointly organizing the field around three guiding questions, it integrates machine‑learning and materials‑science perspectives into a single conceptual framework for autonomous materials discovery. The survey systematically reviews predominantly reactive tasks such as prediction, mining, generation, optimization, and verification, together with cross‑cutting issues in data, knowledge integration, multimodality, and explainability, clarifying both the strengths and limitations of current approaches. Building on this foundation, it characterizes emerging agentic systems and Scientist‑AI workflows that couple hypothesis generation, experiment and simulation planning, and human–AI collaboration, arguing that truly discovery‑oriented AI4MatSci requires aligning all components of the pipeline to real experimental reward signals rather than proxy benchmarks.
Revisiting Fig. 1, the agentic system operating within an open‑ended experimental environment uniquely accesses real‑world outcome signals that can be used for credit assignment at the current decision stage and retroactively propagated to upstream stages spanning pretraining, materials‑specific adaptation, and task design, as detailed in Secs. 2, 3, and 4. From this vantage point, the concluding section elaborates a coherent roadmap that operationalizes this pipeline‑centric view, integrating the insights from the preceding sections into a unified picture of how agentic LLM systems can be aligned with end‑to‑end materials discovery objectives.
@@
Barto 31 observed that “reinforcement learning is memorized search” and noted that “in the beginning, machine learning was reinforcement learning.” From this point of view, general pre-training, domain-specific fine-tuning and part of lifelong RL in the experimental environment serve as a unified memory, contributing to the search component of RL, as shown in Fig. 10. Note that an agentic LLM also relies on the memory function during interacting with the environment, as it stores useful information into the memory after searching.
@@
However, recent evaluations 244 show that frontier LLMs can engage in sophisticated reward hacking in benchmarking environments, indicating that the memory formed by pretraining and fine‑tuning may overfit scoring artifacts rather than reflect robust performance in real experimental settings. To mitigate this, experimental environment rewards should be used to reshape upstream memory, data curation, objectives, and weights, thus aligning learning with outcome‑level goals. It has also been argued 36 that small language models are enough for agentic tasks, as they are “sufficiently powerful, inherently more suitable, and necessarily more economical for many invocations in agentic systems”. To some extent, compressing LLMs into small LMs can be considered as a way to modify the upstream components driven by the credit assignment process mainly based on the final RL reward – discovering novel materials in our case. Ideally, the whole pipeline should be tunable, driven by the final objective.
@@
For MatSci scenarios, we can use the influence function 137 to track down the contributions of pre-training and fine-tuning data to the discovery of novel and useful materials. It is also promising to learn RL rules that are more suitable for MatSci settings, inspired by the success of showing how to discover a state-of-the-art RL rule that outperforms manually designed rules 264. We encourage both the machine learning and materials science communities to view the pipeline as a whole serving the single ultimate goal – discovering novel materials, using which to scrutinize the upstream components and redesign them through the lens of RL.
@@
As indicated in Fig. 10, an agentic LLM can operate on either the virtual world and the real world. It is preferable and practical to first train it on the virtual world and then transfer it and continually train it in the real-world materials science environment. We should be aware that an agent cannot learn beyond the knowledge of a simulator, meaning that it is not realistic to expect that an agent can generalize to the knowledge outside of a simulator designed by human experts and domain data. This indicates the potential performance gap between an agentic LLM trained in the virtual world and the one ideally trained in the real-world environment, which is prohibitively expensive. We also envision parallel memory+search across the virtual and real worlds, similar to the concept of asynchronous thinking proposed by Schilling-Wilhelmi et al. 310.
@@
Current benchmarking methods for computational materials discovery mostly focus on fixed prediction tasks and isolated computational elements, overlooking the inherently iterative, exploratory, and sometimes serendipitous nature of real scientific discovery 236. Even though CrystalGym 134 provides APIs that are familiar to RL researchers, it still remains unclear if the performance can be translated into real-world experimental environments. Inspired by digital twins for materials 177 and co-evolution of agents and the virtual environment 229, where a data interface keeps the digital and physical systems synchronized, we propose to evolve the agent, the virtual world, and all the components in the upstream stages based on the real-world environmental reward. While individual components and their associated views each offer limited perspectives, only a systems-level approach can deliver the complete, integrated understanding required to objectively accelerate materials innovation; this goal underpins the pipeline-centric view of our survey. Here, a modular dynamic AI4MatSci testbed is needed, which can be updated with real-world data more easily.
@@
In terms of additional algorithmic improvement, differently from previous digital-twins for material discovery, we argue for a single shared agent or a population of collective agents with memory 146, 406, 400, 142, 356 for both simulation and real-world experiments, and even upstream pre- and mid-training. This unified view is also aligned with the concept depicted in Fig. 10, where the whole pipeline shares a single memory across all stages, which can be updated by the final real-world reward signal.
@@
Bengio and others 39 warned about the biosecurity risks at the convergence of AI and the life sciences. Similarly, LLM agent-driven materials discovery and autonomous laboratory workflows can accelerate beneficial research but also lower barriers for misuse (e.g., designing materials with hazardous properties or enabling access to dangerous synthesis routes). Thus, it is crucial to develop trustworthy and safe LLM agents for materials science research. We depict the empowerment-plasticity view 4 in Fig. 10. Empowerment is the extent to which an agent can shape or control the set of future states it can perceive, and plasticity is the capacity of an agent to maintain adaptability in the face of significant changes or challenges 4. Interestingly, the plasticity of the agent is equivalent to the empowerment of the environment, and vice-versa 4. On the one hand, this implies that, for the safety concerns, the agent might need to dynamically adjust its plasticity in order to restrain the adversarial empowerment of the environment. On the other hand, this means that to maximize the overall yields of the whole system over time, the environment over which the agent is empowered should often have a high plasticity, e.g., with more controllable components or tools. We argue that AI4MatSci should routinely revisit each module in the workflow and ask whether it can be brought under learning or optimization pressure rather than being treated as a frozen heuristic. Concretely, this means not only training neural networks but also learning data selection and curation policies 59, adaptive pre‑training corpora, retrieval and tool‑calling strategies, experiment proposal policies, and even aspects of the experimental protocol itself, using the ultimate reward signals. Our pipeline‑centric, end‑to‑end perspective thus views the materials discovery loop as a composite system in which as many components as possible are trainable and where signals from successful or failed discoveries can, over time, reshape both models and upstream pipeline choices to better serve the ultimate goal of finding novel, useful, and safe materials. Overall, our proposed unified and pipeline-centric view encourages us to design highly decoupled agents and the corresponding environments to maximize output while minimizing risks.

### v2

@@ 4.4 Pipeline-Centric Perspective
After addressing Q3, we continue our discussion on the end-to-end pipeline-centric perspective, focusing more on the role of agentic LLMs in materials discovery. Yann LeCun has famously argued that “if intelligence is a cake, the bulk of the cake is unsupervised learning, the icing on the cake is supervised learning, and the cherry on the cake is reinforcement learning” (NeurIPS 2016), thereby downplaying the centrality of RL relative to representation learning via large-scale pre-training. In contrast, Silver and Sutton 326 argue that intelligence should fundamentally be grounded in interaction with the environment, thereby elevating the role of RL and other experience-driven approaches to a primary mechanism for building general intelligence. Rather than engaging directly with this broader conceptual debate, we instead focus on the concrete demands of materials science, where discovery is inherently interactive: hypotheses must be tested through simulations and (preferably) experiments, so learning through experience is less optional than it is in purely digital areas.
@@
First, a lot of current work doesn’t use backward credit assignment all the way through the pipeline. By backward credit assignment, we mean linking the results of downstream discoveries—both successes and failures—back to choices made earlier in the process, such as corpus selection, objective design, and model adaptation. This way, those parts can be changed under the same end-to-end pressure. As outlined in Sec. 2 and 3, training signals resulting from agentic successes or failures (such as unsuccessful synthesis attempts, inaccurate property predictions, and breached physical constraints) can theoretically be transmitted backward to prior stages, including pre-training data curation, domain adaptation objectives, and instruction-tuning criteria. This backward signal is necessary for the influence-driven correction discussed in Sec. 2.4, where failures in closed-loop discovery immediately tell us which parts of the pre-training corpus should be made louder, quieter, or eliminated. Without this feedback loop, agentic systems stay separate from the basic representations and assumptions built into their models. This makes it harder for them to change and get better over long periods of time.
@@ 5 Discussion & Conclusion
This survey synthesizes recent advances in AI for materials science through a pipeline‑centric lens that connects corpus curation, pre-training, domain adaptation, and instruction tuning to goal‑conditioned, agentic LLMs operating within open‑ended experimental environments. By jointly organizing the field around three guiding questions, it integrates machine‑learning and materials‑science perspectives into a single conceptual framework for autonomous materials discovery. The survey systematically reviews predominantly reactive tasks such as prediction, mining, generation, optimization, and verification, together with cross‑cutting issues in data, knowledge integration, multimodality, and explainability, clarifying both the strengths and limitations of current approaches. Building on this foundation, it characterizes emerging agentic systems and Scientist‑AI workflows that couple hypothesis generation, experiment and simulation planning, and human–AI collaboration, arguing that truly discovery‑oriented AI4MatSci requires aligning all components of the pipeline to real experimental reward signals rather than proxy benchmarks.
Revisiting Fig. 1, the agentic system operating within an open‑ended experimental environment uniquely accesses real‑world outcome signals that can be used for credit assignment at the current decision stage and retroactively propagated to upstream stages spanning pre-training, materials‑specific adaptation, and task design, as detailed in Secs. 2, 3, and 4. From this vantage point, the concluding section elaborates a coherent roadmap that operationalizes this pipeline‑centric view, integrating the insights from the preceding sections into a unified picture of how agentic LLM systems can be aligned with end‑to‑end materials discovery objectives.
@@
Barto 32 observed that “reinforcement learning is memorized search” and noted that “in the beginning, machine learning was reinforcement learning.” From this point of view, general pre-training, domain-specific fine-tuning and part of lifelong RL in the experimental environment serve as a unified memory, contributing to the search component of RL, as shown in Fig. 8. Note that an agentic LLM also relies on the memory function during interacting with the environment, as it stores useful information into the memory after searching.
@@
However, recent evaluations 246 show that frontier LLMs can engage in sophisticated reward hacking in benchmarking environments, indicating that the memory formed by pre-training and fine‑tuning may overfit scoring artifacts rather than reflect robust performance in real experimental settings. To mitigate this, experimental environment rewards should be used to reshape upstream memory, data curation, objectives, and weights, thus aligning learning with outcome‑level goals. It has also been argued 37 that small language models are enough for agentic tasks, as they are “sufficiently powerful, inherently more suitable, and necessarily more economical for many invocations in agentic systems”. To some extent, compressing LLMs into small LMs can be considered as a way to modify the upstream components driven by the credit assignment process mainly based on the final RL reward – discovering novel materials in our case. Ideally, the whole pipeline should be tunable, driven by the final objective.
@@
For MatSci scenarios, we can use the influence function 138 to track down the contributions of pre-training and fine-tuning data to the discovery of novel and useful materials. It is also promising to learn RL rules that are more suitable for MatSci settings, inspired by the success of showing how to discover a state-of-the-art RL rule that outperforms manually designed rules 266. We encourage both the machine learning and materials science communities to view the pipeline as a whole serving the single ultimate goal – discovering novel materials, using which to scrutinize the upstream components and redesign them through the lens of RL.
@@
As indicated in Fig. 8, an agentic LLM can operate on either the virtual world and the real world. It is preferable and practical to first train it on the virtual world and then transfer it and continually train it in the real-world materials science environment. We should be aware that an agent cannot learn beyond the knowledge of a simulator, meaning that it is not realistic to expect that an agent can generalize to the knowledge outside of a simulator designed by human experts and domain data. This indicates the potential performance gap between an agentic LLM trained in the virtual world and the one ideally trained in the real-world environment, which is prohibitively expensive. A closely related and often underemphasized bottleneck is the limited accessibility of laboratory automation itself. Democratizing lab automation is critical to addressing the persistent data scarcity challenge in materials science. Despite recent progress, autonomous and self-driven laboratories remain expensive, slow to deploy, and heavily dependent on expert knowledge not only in materials science, but also in mechanical engineering, robotics, and software engineering. As a result, adapting an existing automated laboratory to a new scientific objective is still a nontrivial and resource-intensive process, restricting the scale and diversity of real-world feedback available for learning.
We also envision parallel memory+search across the virtual and real worlds, similar to the concept of asynchronous thinking proposed by Schilling-Wilhelmi et al. 312.
@@
Current benchmarking methods for computational materials discovery mostly focus on fixed prediction tasks and isolated computational elements, overlooking the inherently iterative, exploratory, and sometimes serendipitous nature of real scientific discovery 238. Even though CrystalGym 135 provides APIs that are familiar to RL researchers, it still remains unclear if the performance can be translated into real-world experimental environments. Inspired by digital twins for materials 178 and co-evolution of agents and the virtual environment 231, where a data interface keeps the digital and physical systems synchronized, we propose to evolve the agent, the virtual world, and all the components in the upstream stages based on the real-world environmental reward. While individual components and their associated views each offer limited perspectives, only a systems-level approach can deliver the complete, integrated understanding required to objectively accelerate materials innovation; this goal underpins the pipeline-centric view of our survey. Here, a modular dynamic AI4MatSci testbed is needed, which can be updated with real-world data more easily.
@@
In terms of additional algorithmic improvement, differently from previous digital-twins for materials discovery, we argue for a single shared agent or a population of collective agents with memory 147, 409, 403, 143, 359 for both simulation and real-world experiments, and even upstream pre- and mid-training. This unified view is also aligned with the concept depicted in Fig. 8, where the whole pipeline shares a single memory across all stages, which can be updated by the final real-world reward signal.
@@
Bengio and others 40 warned about the biosecurity risks at the convergence of AI and the life sciences. Similarly, LLM agent-driven materials discovery and autonomous laboratory workflows can accelerate beneficial research but also lower barriers for misuse (e.g., designing materials with hazardous properties or enabling access to dangerous synthesis routes). Thus, it is crucial to develop trustworthy and safe LLM agents for materials science research. We depict the empowerment-plasticity view 4 in Fig. 8. Empowerment is the extent to which an agent can shape or control the set of future states it can perceive, and plasticity is the capacity of an agent to maintain adaptability in the face of significant changes or challenges 4. Interestingly, the plasticity of the agent is equivalent to the empowerment of the environment, and vice-versa 4. On the one hand, this implies that, for the safety concerns, the agent might need to dynamically adjust its plasticity in order to restrain the adversarial empowerment of the environment. On the other hand, this means that to maximize the overall yields of the whole system over time, the environment over which the agent is empowered should often have a high plasticity, e.g., with more controllable components or tools. We argue that AI4MatSci should routinely revisit each module in the workflow and ask whether it can be brought under learning or optimization pressure rather than being treated as a frozen heuristic. Concretely, this means not only training neural networks but also learning data selection and curation policies 60, adaptive pre‑training corpora, retrieval and tool‑calling strategies, experiment proposal policies, and even aspects of the experimental protocol itself, using the ultimate reward signals. Our pipeline‑centric, end‑to‑end perspective thus views the materials discovery loop as a composite system in which as many components as possible are trainable and where signals from successful or failed discoveries can, over time, reshape both models and upstream pipeline choices to better serve the ultimate goal of finding novel, useful, and safe materials. Overall, our proposed unified and pipeline-centric view encourages us to design highly decoupled agents and the corresponding environments to maximize output while minimizing risks.